# How Does A Text Preprocessing Pipeline Affect Syntactic Ontology Matching?

### v8

@@ 5.4. Post Hoc LLM-Based Repair
    We choose 10 LLMs from 4 different families. These include two OpenAI GPT models (OpenAI, [n. d.]) (gpt-4o and gpt-4o-mini), two Anthropic Claude models (Anthropic, [n. d.])(claude-3-sonnet and claude-3-haiku), and two versions of Meta Llama (Meta, [n. d.]) (llama-3-8b and llama-3.1-8b), two versions of Alibaba Qwen (Alibaba Qwen Team, [n. d.]) (qwen-2-7b and qwen-2.5-7b), Google Gemma 2 (Google Gemma Team and Google DeepMind, [n. d.]), and ChatGLM 4 (Zeng et al., 2024). We use LangChain (LangChain, Inc., [n. d.]) to build chats with OpenAI GPT and Anthropic Claude models, and Ollama (Ollama, [n. d.]) to access open-source models. Table 3 lists the details of selected LLMs. All model temperatures are set to 0 to minimise the random results generated by LLMs.
@@
    Table 3. Details of selected LLMs (retrieved from (Qiang et al., 2023)). API-accessed models are shown as triangles ($\blacktriangle$) and open-source models as circles ($\bullet$).
@@
    We select 4 different LLM prompt strategies. (a) Basic: The simplest prompt template to compare two entities using “Is Entity1 equivalent to Entity2? Answer yes or no.” (b) Few-Shot: Adding an example in the prompt. We use the example “Example: Hair_root is equivalent to Hair_Root.” (c) Self-Reflection: Before LLMs answer the question, ask LLMs to rethink the output and give a short explanation. We add a sentence “Write a short explanation.” to invoke this function. (d) Few-Shot + Self-Reflection: A combination of (b) and (c). Table 4 lists the details of the selected prompt strategies and their templates.
@@
    Table 4. Selected prompt templates. The few-shot parts are coloured blue, and the self-reflection parts are orange.
@@
    No. Prompt Strategy Prompt Template PT1 Basic Is Entity1 equivalent to Entity2? Answer yes or no. PT2 Few-Shot Example: Hair_root is equivalent to Hair_Root. Is Entity1 equivalent to Entity2? Answer yes or no. PT3 Self-Reflection Is Entity1 equivalent to Entity2? Answer yes or no. Write a short explanation. PT4 Few-Shot Example: Hair_root is equivalent to Hair_Root. + Is Entity1 equivalent to Entity2? Answer yes or no. Self-Reflection Write a short explanation.
  4. (4)
@@
    We compare LLMs with the classical text preprocessing pipeline in detecting TPs and FPs, which are two main contributors to precision and recall in matching performance. The true mappings (TPs) are retrieved from the reference (R) where the mappings can be found by syntactic OM. The false mappings (FPs) are retrieved from the alignment (A) generated by the classical text preprocessing pipeline but exclude TPs above.
  5. (5)
    We report our results over a single run. There could be slight differences in the results over multiple trials, but these differences are not significant with respect to the ranking and trend of different LLMs and prompt templates.
@@
Figure 10(l) compares LLMs with the classical text preprocessing pipeline in detecting TPs. LLMs and the text preprocessing pipeline are expected to answer “yes” for TPs. The discovery rate is defined as the number of “yes” answers compared to the total number of TPs. For LLMs, we directly prompt LLMs to determine whether two entities are identical or not. For the text preprocessing pipeline, two entities are considered to be matched if they have the same results after the text preprocessing. Here we use only Tokenisation (T) and Normalisation (N) to preprocess the text because Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are shown to be less effective in our analysis in Section 4. Although some API-accessed commercial LLMs have shown the capability to detect TPs without a text preprocessing step, this behaviour varies across different LLMs and prompt templates. For different LLMs, open-source LLMs perform better than commercial API-accessed LLMs. It could be problematic for some commercial API-accessed LLMs (e.g. claude-3-haiku) to find two identical entities if text preprocessing is not applied prior to the comparison. They may consider entities with different formatting that are used in different contexts and therefore have different meanings. For different prompt templates, few-shot (PT2) significantly improves correctness, but self-reflection (PT3) may even hamper performance. The combination of few-shot and self-reflection (PT4) may even perform worse than the simple basic prompt (PT1). We believe that this phenomenon is mainly attributed to “overthinking”, a type of LLM hallucination in which LLMs could over-analyse user input and generate unnecessarily complex details (Chen et al., 2025). We also need to note here that the example generated in the few-shot prompt is based on our observation of the existing TPs. Finding a good example in real-world applications is not easy because ontology entities can have different formatting by nature. On the other hand, using the classical text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.
Figure 10. Discovery rate of TPs using LLMs on the Largebio Track.
[…]
Discovery rate of TPs using LLMs on the Largebio Track.
Figure 11. The discovery rate of FPs using LLMs on the Largebio Track.
[…]
Discovery rate of FPs using LLMs on the Largebio Track.
@@
Figure 11(l) compares LLMs with the classical text preprocessing pipeline in detecting FPs. LLMs and the text preprocessing pipeline are expected to answer “no” for FPs. The discovery rate is defined as the number of “no” answers compared to the total number of FPs. Specifically, these FPs evaluated are non-trivial mappings that cannot be detected by the classical text preprocessing method. Here we use LLMs to detect FPs generated by the Lancaster Stemmer (SL). SL has shown a lower precision compared to other text preprocessing methods, which is caused by the method producing a significant number of FPs. LLMs have shown a strong capability to discover the FPs produced by the classical text preprocessing pipeline. For different models, API-accessed commercial LLMs are better than open-source LLMs and reach a higher discovery rate of more than 80%. For different prompt templates, both few-shot (PT2), self-reflection (PT3), and a combination of the above (PT4) are better than a simple basic prompt (PT4). While mapping repair is usually non-trivial and requires domain expertise, LLMs have shown potential for performing a high-quality post hoc correction without human intervention.
For syntactic OM, we can see that LLMs are not good at detecting true mappings, but they are good at detecting false mappings. Therefore, we propose our generic post hoc LLM-based repair with
  1. (1)
@@
    Employ the classical text preprocessing pipeline via function calling (a.k.a. tool calling) to detect TPs. This means that the LLMs act as an interface, and the mapping judgement is based on the programme logic defined in the text preprocessing method. This provides stable results for text preprocessing when finding FPs. LLM prompting is not suitable for this task, even with complex prompting strategies.
  2. (2)
    Use LLM prompting to detect FPs. The prompt template can be simple like “Is A equivalent to B?”, as we cannot see significant improvement using complex prompting strategies (except Meta Llama models).
5.5. Evaluation
@@
We apply our pipeline repair approaches on the Phase 2 text preprocessing method Lancaster Stemmer (SL), and evaluate their performance on the three large datasets in the Largebio Track: WHOLE-FMA-NCI, WHOLE-FMA-SNOMED, and WHOLE-SNOMED-NCI.
Figure 12(c) compares the matching performance with/without using the pipeline approaches: ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of the above (CR). We can see that ad hoc logic-based repair (AR) and post hoc LLM-based repair (PR) can significantly improve precision, with only a slight decrease in recall, but the overall F1 score also increases. Although there is no restriction on the use of both ad hoc and post hoc repairs, the combination does not improve precision, recall, or F1 score. For syntactic OM, its main purpose is to provide “seed mappings” for lexical and semantic matching. These mappings are meant to be precise and accurate. Both of our pipeline repair approaches have shown high precision and are the best tool suites.
(a) WHOLE-FMA-NCI
(b) WHOLE-FMA-SNOMED
(c) WHOLE-SNOMED-NCI
Figure 12. Evaluation of pipeline repair approaches: Ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of above (CR). The symbol $\varnothing$ represents the baseline without text preprocessing repair.
[…]
Evaluation of pipeline repair approaches.
Figure 13 evaluates the scalability of ad hoc logic-based repair (AR). The x-axis indicates the number of entities in each alignment, and the y-axis shows the number of reserved words per 100 entities. For all the alignments we tested, the value of reserved words per 100 entities is less than 10. With the number of entities increasing (i.e. the ontology size increases), the number of reserved words per 100 entities decreases. This may be because reserved words such as“has”, “is”, and “a” are commonly used in many ontologies and in multiple distinct entities within those ontologies. Conventionally there are a limited number of such words. Therefore the reserved word feature of our approach is scalable in large-scale OM.
Figure 13. Reserved words per 100 entities vs the number of entities in each alignment.
[…]
Comparison of reserved words per 100 entities (@100) with the number of entities in each alignment.
@@
Figure 14(c) compares post hoc LLM-based repair (PR) with two alternative repair approaches. The results show that PR achieves almost the same precision with text preprocessing only repair (TOR) and LLM only repair (LOR), but its recall and F1 score are better than the other two alternatives. For syntactic OM, this means that our approach could detect more ”seed mappings”.
(a) WHOLE-FMA-NCI
(b) WHOLE-FMA-SNOMED
(c) WHOLE-SNOMED-NCI
@@
Figure 14. Comparison of post hoc LLM-based repair (PR) with two alternative repair approaches: text preprocessing only repair (TOR) and LLM only repair (LOR).
[…]
Comparison of post hoc LLM-based repair (PR) with two alternative repair approaches.
6. Limitations
  1. (1)
    The references contained in the OAEI tracks are not perfect gold standards and some need further development. Therefore at present 100% accuracy is not necessarily desirable.
  2. (2)
@@ 6. Limitations
    In this study we deal only with equivalence mappings between classes and properties. Subsumption mappings are not considered. Finding widely used baseline data with ground truth matches for subsumptions is empirically difficult. Only a few OAEI tracks contain subsumption mappings, and they are not enough to robustly demonstrate the text preprocessing used in subsumption mappings.
  3. (3)
    This work has not been included in the official OAEI campaign because of its intense focus on text preprocessing in syntactic OM.
7. Conclusion
@@
In this paper, we conduct a comprehensive study on the effect of the text preprocessing pipeline on syntactic OM. 8 OAEI tracks with 49 distinct alignments are evaluated. Despite the importance of text preprocessing in syntactic OM, our experimental results indicate that the text preprocessing pipeline is currently ill-equipped. (1) We find that Phase 1 text preprocessing methods (Tokenisation and Normalisation) help with both matching completeness (i.e. recall) and correctness (i.e. precision). Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are less effective. They can improve matching completeness (i.e. recall), but matching correctness (i.e. precision) is relatively low. (2) We propose two novel pipeline repair methods to repair the less effective Phase 2 text preprocessings. Both approaches outperform the traditional text preprocessing pipeline without repair, in particular, the matching correctness (i.e. precision) and overall matching performance (i.e. F1 score).
@@
While the classical text preprocessing pipeline in syntactic OM is usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
  1. (1)
@@
    Our study shows (a) whether to use, or not use, and (b) how to use these text preprocessing methods. Such experimental results will benefit the selection of appropriate text preprocessing methods for syntactic OM. For large-scale OM, it can significantly reduce unnecessary trial costs.
  2. (2)
@@
    Our pipeline repair approaches can be used to repair the less effective Phase 2 text preprocessing methods. Its broader value is showing how to maximise true mappings and minimise false mappings throughout the text preprocessing pipeline. From a statistical perspective, it is a local optimisation for syntactic matching that will benefit the global optimisation for OM, as syntactic matching provides “anchor mappings” for lexical and semantic matching.
  3. (3)
@@
    We showcase the limitations and opportunities for integrating the text preprocessing pipeline with LLMs in syntactic OM. We find that LLMs are good tools to find false mappings, but their outputs for finding true mappings are non-deterministic. The best practice is to integrate the classical text preprocessing pipelines with the LLMs. Using function calling to invoke the classical text preprocessing pipeline within LLMs, our approach could produce more stable and robust “seed mappings” than traditional text preprocessing methods and purely prompt-based approaches.
  4. (4)
@@
    Our study also demonstrates that the nature of text preprocessing used in NLP applications can be context-based. Not all pipeline methods are effective, and some of them may even hamper the performance of downstream tasks. Modifications are required before applying text preprocessing to specific NLP tasks.
@@
Our future work will focus on handling class axioms and complex relationships to evaluate the text preprocessing pipeline. We will also study the pipeline and pipeline repair approach working with both knowledge-based OM systems and LLM-based OM systems.
Acknowledgements.

### v10

@@ V-D Post Hoc LLM-Based Repair
We choose 10 LLMs from 4 different families. These include two OpenAI GPT models [50] (gpt-4o and gpt-4o-mini), two Anthropic Claude models [51] (claude-3-sonnet and claude-3-haiku), and two versions of Meta Llama [52] (llama-3-8b and llama-3.1-8b), two versions of Alibaba Qwen [53] (qwen-2-7b and qwen-2.5-7b), Google Gemma 2 [54], and ChatGLM 4 [55]. We use LangChain [56] to build chats with OpenAI GPT and Anthropic Claude models, and Ollama [57] to access open-source models. Table II lists the details of selected LLMs (retrieved from [14]). All model temperatures are set to 0 to minimise the random results generated by LLMs.
@@
TABLE II: Details of selected LLMs. API-accessed models are shown as triangles ($\blacktriangle$) and open-source models as circles ($\bullet$).
@@
We select 4 different LLM prompt strategies. (a) Basic: The simplest prompt template to compare two entities using “Is Entity1 equivalent to Entity2? Answer yes or no.” (b) Few-Shot: Adding an example in the prompt. We use the example “Example: Hair_root is equivalent to Hair_Root.” (c) Self-Reflection: Before LLMs answer the question, ask LLMs to rethink the output and give a short explanation. We add a sentence “Write a short explanation.” to invoke this function. (d) Few-Shot + Self-Reflection: A combination of (b) and (c). Table III lists the details of the selected prompt strategies and their templates.
@@
TABLE III: Selected prompt templates. The few-shot parts are coloured blue, and the self-reflection parts are orange.
@@
No. Prompt Strategy Prompt Template PT1 Zero-Shot Is Entity1 equivalent to Entity2? Answer yes or no. PT2 Few-Shot Example: Hair_root is equivalent to Hair_Root. Is Entity1 equivalent to Entity2? Answer yes or no. PT3 Self-Reflection Is Entity1 equivalent to Entity2? Answer yes or no. Write a short explanation. PT4 Few-Shot Example: Hair_root is equivalent to Hair_Root. + Is Entity1 equivalent to Entity2? Answer yes or no. Self-Reflection Write a short explanation.
@@
We experiment with the Largebio Track that contains pairwise alignments between FMA, NCI, and SNOMED. We choose three datasets: WHOLE-FMA-NCI, WHOLE-FMA-SNOMED, and WHOLE-SNOMED-NCI. We compare LLMs with the classical text preprocessing pipeline in detecting TPs and FPs, which are two main contributors to precision and recall in matching performance. The true mappings (TPs) are retrieved from the reference (R) where the mappings can be found by OM. The false mappings (FPs) are retrieved from the alignment (A) generated by the classical text preprocessing pipeline but exclude TPs above. In this study, we report our results on a single run. There may be slight differences in the results across multiple trials, but these differences are not significant in terms of the ranking and trend of different LLMs and prompt templates.
@@
Fig. 13 compares LLMs with the classical text preprocessing pipeline in detecting TPs. LLMs and the text preprocessing pipeline are expected to answer “yes” for TPs. The discovery rate is defined as the number of “yes” answers compared to the total number of TPs. For LLMs, we directly prompt LLMs to determine whether two entities are identical or not. For the text preprocessing pipeline, two entities are considered to be matched if they have the same results after the text preprocessing. Here we use only Tokenisation (T) and Normalisation (N) to preprocess the text because Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are shown to be less effective in our analysis in Section IV. Although some API-accessed commercial LLMs have shown the capability to detect TPs without a text preprocessing step, this behaviour varies across different LLMs and prompt templates. For different LLMs, open-source LLMs perform better than commercial API-accessed LLMs. It could be problematic for some commercial API-accessed LLMs (e.g. claude-3-haiku) to find two identical entities if text preprocessing is not applied prior to the comparison. They may consider entities with different formatting that are used in different contexts and therefore have different meanings. For different prompt templates, few-shot (PT2) significantly improves correctness, but self-reflection (PT3) may even hamper performance. The combination of few-shot and self-reflection (PT4) may even perform worse than the simple basic prompt (PT1). We believe that this phenomenon is mainly attributed to “overthinking”, a type of LLM hallucination in which LLMs could over-analyse user input and generate unnecessarily complex details [58]. We also need to note here that the example generated in the few-shot prompt is based on our observation of the existing TPs. Finding a good example in real-world applications is not easy because ontology entities can have different formatting by nature. On the other hand, using the classical text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.
(a) PT1: WHOLE-FMA-NCI
(b) PT1: WHOLE-FMA-SNOMED
(c) PT1: WHOLE-SNOMED-NCI
(d) PT2: WHOLE-FMA-NCI
(e) PT2: WHOLE-FMA-SNOMED
(f) PT2: WHOLE-SNOMED-NCI
(g) PT3: WHOLE-FMA-NCI
(h) PT3: WHOLE-FMA-SNOMED
(i) PT3: WHOLE-SNOMED-NCI
(j) PT4: WHOLE-FMA-NCI
(k) PT4: WHOLE-FMA-SNOMED
(l) PT4: WHOLE-SNOMED-NCI
Figure 13: Discovery rate of TPs using LLMs on the Largebio Track.
@@
Fig. 14 compares LLMs with the classical text preprocessing pipeline in detecting FPs. LLMs and the text preprocessing pipeline are expected to answer “no” for FPs. The discovery rate is defined as the number of “no” answers compared to the total number of FPs. Specifically, these FPs evaluated are non-trivial mappings that cannot be detected by the classical text preprocessing method. Here we use LLMs to detect FPs generated by the Lancaster Stemmer (SL). SL has shown a lower precision compared to other text preprocessing methods, which is caused by the method producing a significant number of FPs. LLMs have shown a strong capability to discover the FPs produced by the classical text preprocessing pipeline. For different models, API-accessed commercial LLMs are better than open-source LLMs and reach a higher discovery rate of more than 80%. For different prompt templates, both few-shot (PT2), self-reflection (PT3), and a combination of the above (PT4) are better than a simple basic prompt (PT4). While mapping repair is usually non-trivial and requires domain expertise, LLMs have shown potential for performing a high-quality post hoc correction without human intervention.
(a) PT1: WHOLE-FMA-NCI
(b) PT1: WHOLE-FMA-SNOMED
(c) PT1: WHOLE-SNOMED-NCI
(d) PT2: WHOLE-FMA-NCI
(e) PT2: WHOLE-FMA-SNOMED
(f) PT2: WHOLE-SNOMED-NCI
(g) PT3: WHOLE-FMA-NCI
(h) PT3: WHOLE-FMA-SNOMED
(i) PT3: WHOLE-SNOMED-NCI
(j) PT4: WHOLE-FMA-NCI
(k) PT4: WHOLE-FMA-SNOMED
(l) PT4: WHOLE-SNOMED-NCI
Figure 14: The discovery rate of FPs using LLMs on the Largebio Track.
@@
For OM, we can see that LLMs are not good at detecting true mappings, but they are good at detecting false mappings. Therefore, we propose our post hoc LLM-based repair with two steps. (1) Given an alignment, employ the classical text preprocessing pipeline on each mapping via an LLM function calling (a.k.a. tool calling) to detect TPs. This means that the LLMs act as an interface, and the mapping judgement is based on the program logic defined in the text preprocessing method. This provides stable results for text preprocessing when finding FPs. LLM prompting is not suitable for this task, even with complex prompting strategies. (2) Given an alignment, use LLM prompting on each mapping to detect FPs. The prompt template can be simple like “Is A equivalent to B?”, as we cannot see significant improvement using complex prompting strategies (except for Meta Llama models).
@@
We apply our pipeline repair approaches on the Phase 2 text preprocessing method Lancaster Stemmer (SL). Due to the superior performance and superior cost of the best LLMs, we evaluate only on the three largest datasets in the Largebio Track: WHOLE-FMA-NCI, WHOLE-FMA-SNOMED, and WHOLE-SNOMED-NCI. With this choice, we expect the performance increase may justify the cost of the repair step.
@@
Fig. 15 compares post hoc LLM-based repair (PR) with two alternatives: text-preprocessing-only repair (TOR) and LLM-only repair (LOR). The results show that PR achieves almost the same precision as TOR and LOR, but its recall and F1 score are better than the other two alternatives. For OM, this means that our approach could detect more ”seed mappings”.
Fig. 16 compares the matching performance with/without using the pipeline repair approaches: ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of the above (CR). We can see that AR and PR can significantly improve precision, with only a slight decrease in recall, but the overall F1 score also increases. For text preprocessing used in OM, its main purpose is to provide “seed mappings” for lexical and semantic matching. These mappings are meant to be precise and accurate. Both of our pipeline repair approaches shown with high precision are the best tool suites for OM. Although there is no restriction on the use of both ad hoc and post hoc repairs, the combination CR does not improve precision, recall, or F1 score.
@@
Figure 15: Comparison of post hoc LLM-based repair (PR) with two alternatives: text-preprocessing-only repair (TOR) and LLM-only repair (LOR).
Figure 16: Evaluation of pipeline repair approaches: Ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of above (CR). The symbol $\varnothing$ represents the baseline without text preprocessing repair.
VI Limitations
@@
The references contained in the OAEI tracks are not perfect gold standards and some need further development. Therefore at present 100% accuracy is not necessarily desirable. In this study, we deal only with equivalence mappings between classes and properties. Subsumption mappings are not considered. Finding widely used baseline data with ground truth matches for subsumptions is empirically difficult. Only a few OAEI tracks contain subsumption mappings, and they are not enough to robustly demonstrate the text preprocessing used in subsumption mappings.
VII Conclusion
@@
In this paper, we conduct a comprehensive study on the effect of the text preprocessing pipeline on OM. 8 OAEI tracks with 49 distinct alignments are evaluated. Despite the importance of text preprocessing in OM, our experimental results indicate that the text preprocessing pipeline is currently ill-equipped. We find that Phase 1 text preprocessing methods (Tokenisation and Normalisation) help with both matching completeness (i.e. recall) and correctness (i.e. precision). Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are less effective. They can improve matching completeness (i.e. recall), but matching correctness (i.e. precision) is relatively low. We propose two novel pipeline repair methods to repair the less effective Phase 2 text preprocessings. Both approaches outperform the traditional text preprocessing pipeline without repair, in particular, the matching correctness (i.e. precision) and overall matching performance (i.e. F1 score). We argue that the traditional text preprocessing pipeline is not obsolete in the era of LLMs. It is an integral part of the LLMs used in OM tasks. Integrating the classical text preprocessing pipeline can significantly improve the robustness of purely prompt-based approaches.
@@
While the text preprocessing pipeline in OM is usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
@@
  • Our study shows (a) whether to use, or not use, and (b) how to use these text preprocessing methods. Such experimental results will benefit the selection of appropriate text preprocessing methods for OM. For large-scale OM, it can significantly reduce unnecessary trial costs.
@@
  • Our pipeline repair approaches can be used to repair the less effective Phase 2 text preprocessing methods. Its broader value is showing how to maximise true mappings and minimise false mappings throughout the text preprocessing pipeline. From a statistical perspective, it is a local optimisation for syntactic matching that will benefit the global optimisation for OM.
@@
  • We showcase the opportunities for integrating the text preprocessing pipeline with LLMs in OM. We find that LLMs are good tools to find false mappings, but their outputs for finding true mappings are non-deterministic. The best practice is to integrate the classical text preprocessing pipelines with the LLMs. Using function calling to invoke the classical text preprocessing pipeline within LLMs, our approach could produce more stable and robust “seed mappings” than traditional text preprocessing methods and purely prompt-based approaches.
@@
  • Our study demonstrates that the nature of text preprocessing used in downstream applications can be context-based. Not all methods are effective, and some of them may even hamper the performance of downstream tasks. Modifications are required before applying text preprocessing methods to specific tasks.
@@
Our future work will focus on handling class axioms and complex relationships to evaluate the text preprocessing pipeline. We will also study the pipeline and pipeline repair approach working with both knowledge-based OM systems and LLM-based OM systems. This work has been included in the Agent-OM entry to the 2025 OAEI campaign.
Artifacts
The source code, data, and/or other artifacts have been made available at https://github.com/qzc438-research/ontology-nlp. The OAEI datasets are available from the OAEI Matching EvaLuation Toolkit (MELT) at https://dwslab.github.io/melt/track-repository (retrieved January 1, 2025). The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
Acknowledgment